OpenStudy (anonymous):

If an object is thrown straight upward with an initial sped of 8m/s and takes 3 seconds to strike the ground, from what height was the object thrown?

OpenStudy (anonymous):

"takes three seconds to strike the ground" means from being tossed or from the top of its trajectory? 8 m/s will come to a stop when 8 m/s- g t = 0 t = (8/9.8) s the falling part will be distance s = (1/2) (g) t^2, with t =3 or perhaps 3 -(8/9)s, depending on interpretation.